OpenStudy (anonymous):

find the number of possible positive real zeros of f(x)=8x^4-72x^3+144x^2

OpenStudy (paxpolaris):

\[\large f(x)=8x^4-72x^3+144x^2 \] is this it?

OpenStudy (anonymous):

yes

OpenStudy (paxpolaris):

\[\large f(x)=8x^2\left( x^2-9x+18 \right)\]

OpenStudy (anonymous):

wait so what do i do?? im so confuseddd

OpenStudy (paxpolaris):

we can see that f(x)=0 has 1 repeated at x=0. next we can check if there are any positive solutions to x^2-9x+18=0

OpenStudy (paxpolaris):

\[\large f(x)=8x^2\left(x-6 \right)\left( x-3 \right)\] ^^^ that's the factored form of f(x) ... correct ?

OpenStudy (paxpolaris):

so \(f(x)=0\) when \(x=0, 6 ,\) or \(3\)

OpenStudy (paxpolaris):

so there are \(\Large 2\) positive, real zeroes : 3 and 6.

OpenStudy (anonymous):

ohhhh okay, so whenever there is a 0 with problems like this it doesnt count as a positive real zero right ?

OpenStudy (paxpolaris):

0 is real ... but not positive.

OpenStudy (anonymous):

oh okay!! thank you so much !

OpenStudy (paxpolaris):

but .... the question isn't asking you actually find the zeroes ... just how many are positive and real. there is a way to do that... even when f(x) can't be so easily factored: www.purplemath.com/modules/drofsign.htm